OpenStudy (anonymous):

Please help! What is the simplest form of the expression? √6x^8y^9 / √5x^2y^4 A. x^3y^2 √30y / 5 B. √30x^10y^13 / 5x^2y^4 C. 5x^3y^2 √30y D. none of these

OpenStudy (anonymous):

Clarification: is this what your expression actually looks like? \[\frac{ \sqrt{6x ^{8}y ^{9}} }{ \sqrt{5x ^{2}y ^{4}} }\]

OpenStudy (anonymous):

Ye. @Cardinal_Carlo

OpenStudy (anonymous):

i think c im not sure yet

OpenStudy (anonymous):

Okay, so in that case, we first merge both rooted expressions: \[\sqrt{\frac{ 6x ^{8}y ^{9} }{ 5x ^{2}y ^{4} }}\] then we simplify: \[\sqrt{(\frac{ 6 }{ 5 }) x ^{6}y ^{5}}\] \[x ^{3}y ^{2}\sqrt{(\frac{ 6 }{ 5 })y}\] And since the answer is not up there, it's D none of these.

OpenStudy (anonymous):

Please help? @Cardinal_Carlo

OpenStudy (anonymous):

Nevermind! Thank you so much :) :)

OpenStudy (anonymous):

You're welcome